40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 8 01:58:53 2014, Jenne, Update, LSC, Playing with PRMI + 2 arms 
    Reply  Tue Apr 8 10:44:53 2014, Koji, Update, LSC, Playing with PRMI + 2 arms 
Message ID: 9787     Entry time: Tue Apr 8 01:58:53 2014     Reply to this: 9788
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Playing with PRMI + 2 arms 

I played around with the PRMI + 2 arms situation again this evening.  (I'm not ready to call it "PRFPMI" until we're at least partially using IR signals for CARM and DARM control).

I'm still a little bothered by the fact that we lose almost all light in the PRC when we're reducing the CARM offset.  I'm not sure where the light is going, but it's not circulating in the PRC, since I see the POP camera get very dark.  I can bring back light by changing either the PRCL offset, the MICH offset, or to a lesser extent (or maybe I'm not going far enough in offset) the DARM offset.

Tonight, I was using ALS to push on the ETMs in DARM/CARM mode (I didn't push on MC2 today, since it was being finicky and I had a hard time locking CARM with the MC as the actuator today). 

For the PRMI, I was using REFL 33 I&Q.  PRCL gain was -0.04, MICH gain was +0.8.  REFL 33I varied between +2 and +1.2 (smaller gain necessary as CARM offset was reduced, but it's easier to acquire at large CARM offset with larger gain), and REFL 33Q was +1.0.   PRCL has the usual FM 2,3,6,9 triggered, but MICH only had FM 2 triggered.  The others (particularly FM6) make MICH lose lock.  PRCL ASC was engaged, with the PRM oplev off.

Most of what I was trying tonight was to reduce the CARM offset, and then adjust some other offset (MICH, PRCL or DARM) to maximize POP DC.  It's possible that the POP 110 and 22 diodes are changing their optimal demod phases as the optical plant changes, but POPDC is just DC. 

I wasn't able to get the CARM offset below about 40 counts.  At some point I had both CARM and DARM offsets at 50 counts, and had IR resonance in the Xarm, but no resonance in the Yarm.  I guess this is just part of having a simultaneous CARM and DARM offset.

EDIT:  If I leave the CARM offset at 0, and use a large DARM offset (100 counts), I can acquire PRMI lock, but I run into the same problems as I reduce the DARM offset - the POP power decreases, and I lose lock around 45 counts.


Side note: Earlier today I redid the POP 110 and 22 phases.  POP110 was -61 deg, and is not -101 deg.  POP22 was +164 deg, and is now -105 deg.  I'm not sure why they needed such radical changes.  When sideband locked on REFL33, POP110 had an average DC level of 580 cts, while POP22 had a value of 290 cts.

ELOG V3.1.3-